OpenStudy (jh99):

How would you simply explain to someone how to solve cot x/2 = 1

ganeshie8 (ganeshie8):

Depends. If that someone knew that \(1 = \cot(\pi/4 + n\pi)\), he would find below steps simple enough : \(\cot(x/2) = 1 = \cot (\pi/4+n\pi)\) \(x/2 = \pi/4 + n\pi\) \(x = \pi/2 + 2n\pi\)

OpenStudy (jh99):

what about quadratic equations? 2 sin^2 x - sin x - 1 = 0